Solve the system graphically or algebraically.
x-y=3
x-y^2=1

Answer by stanbon(75887)

Solve the system graphically or algebraically.
x-y=3
x-y^2=1
----------------
Solve for "x":
x = y+3
-----
Substitute for "x" and solve for "y":
y+3 - y^2 = 1
----
y^2 - y -2 = 0
(y-2)(y+1) = 0
---
y = 2 or y = -1
---
If y = 2, x = y+3 = 5
If y = -1, x = y+3 = 2
---------------------------
Cheers,
Stan H.
======================